OPTICAL-OPTICAL DOUBLE RESONANCE SPECTROSCOPY OF SrOH: THE C̃ ^{2}Π(000) − Ã ^{2}Π(000) AND THE B̃′ ^{2}Σ^{+}(000) − Ã^{2}Π(000) TRANSITIONS

\underline{J.-G. WANG}, P. M. SHERIDAN, Department of Chemistry, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada; M. J. DICK, Department of Physics, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada; S. YU and P. F. BERNATH, Department of Chemistry, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada.

The C̃ ^{2}Π(000) − Ã ^{2}Π(000) transition of SrOH has been rotationally analyzed using optical-optical double resonance (OODR) spectroscopy. SrOH was synthesized in a Broida-type oven by the reaction of H_{2}O_{2} and strontium vapor. The OODR spectrum was measured using a broad band dye laser (∼1 cm^{−1} bandwidth) and a single mode Ti:Sapphire laser as the pump (Ã ^{2}Π − X̃ ^{2}Σ^{+}) and probe (C̃ ^{2}Π − Ã ^{2}Π) lasers, respectively. Rotational and fine structure parameters have been determined through a combined least-squares fit with the Ã ^{2}Π − X̃ ^{2}Σ^{+} optical transition data and the millimeter-wave pure rotational data of the X̃ ^{2}Σ^{+} state. A significant decrease in the spin-orbit constant from the Ã^{2}Π to the C̃ ^{2}Π state has been observed and can be rationalized by the different atomic orbital character of the two states. The Λ-doubling constants of the C̃ ^{2}Π state have been examined using the pure procession model and suggest that this state forms a unique perturber pair with the nearby D̃ ^{2}Σ^{+} state. In addition, the rotationally-resolved spectrum of the B̃′ ^{2}Σ^{+}(000) − Ã^{2}Π(000) transition has been recorded using the OODR technique. Assignment and analysis of this spectrum is in progress and preliminary results will also be presented.